OpenStudy (anonymous):

1. 3(2x – 1) = 15 2. 6(x + 1) = 12 3. 2(4x + 3) = 22 solve for x

OpenStudy (callisto):

3(2x – 1) = 15 (2x – 1) = 5 2x=6 x=3

OpenStudy (callisto):

6(x + 1) = 12 x+1 = 2 x=1

OpenStudy (anonymous):

Although Callisto's method is correct, I'd suggest practising using the distributive property, as it seems to be what the question is looking for.

OpenStudy (callisto):

2(4x + 3) = 22 4x+3 = 11 4x =8 x=2

OpenStudy (anonymous):

Thanks guys

OpenStudy (anonymous):

1. 3 (2 x - 1) = 15 2. 6 (x + 1) = 12 3. 2 (4 x + 3) = 22